MODS1 Commissioning Run 5: UTC 2011 February 14

Observer(s): R. Pogge & J. Antognini (OSU), M. Pedani (LBTO)
Telescope Operator: G. Bechetti (LBTO)
AO Support: Remote from Tucson as needed
Telescope Support: M. Midkiff (LBTO)
Instrument Support: D. Officer (LBTO)

Summary

First night of the February Commissioning Run.

Heavy high clouds moving in late afternoon. Pointing and collimation startup was a little ragged, but we weren't feeling rushed and the passing thick cirrus added to the problems.

We lost about 7 hours to heavy clouds, by early morning, however, we managed to finally catch a break and got 2 hours on-sky. Tested DETXY offsets for putting targets onto a long slit at PA=0 and PA=45, measuring and confirming the sign and scale for DETXY coordinates.

Not a great start, but at least we got on sky after a hiatus of 14 straight nights since Nov 19.

Preparations

Got MODS1 fired up w/o problems - I elected to do a full restart of the IEBs and IUB to run through the sequence. Logs will show multiple fault messages as I exercised fault condition checking. Startup went OK, got notes for the startup script. HEBs were on from before and behaving well. IMCS locked on right away at Zenith. Test bias images looked good, transfer went w/o problems.

Code Updates:
  • MODS observing script engines, transitioning from engineering v0.x beta to observing v1.0
  • modsUI v1.9.11, build of 2011 Jan 26 - minor tweaks after the January run
  • lbttcs v2.6.6 - minor tweaks from the January run.

Details

All times are UTC unless otherwise noted.

0100 - Nice sunset, but mostly cloudy in the west/southwest with high clouds. East is more clear, but still has thin stuff. Winds are 3-5m/s out of the SW, Temp 5°C, and dry with 4% RH.

Startup

0200 - Going to our first 4th mag pointing star through cirrus. Geno having some problems with launching the preset from IRAF (it is complaining about trying to create a log and failing to make a directory for it).

Instead I selected the star (M5_1133) from the BrightStars catalog and sent the preset from the MODS control panel. Executing w/o problems so far.

0230 - waiting for clouds to thin enough to get a clue about the star. We're a bit off... Tweaking pointing terms...
   IE = -2
   CA = +15
Confirmatory acquire preset to ACT0344. Very close, tweaking up another bit
   IE = +5
   CA = +20

0240 - On-axis active preset to BS9116 to tweak collimation. Primary Mirror went into its limit during the preset, input from the collimation lookup table?

When Geno tried to get it off the limits, the image went way out of focus. Resent it as an acquire preset, not having any luck.

0300 - Call is in to John Hill for assistance with collimation...

0320 - Changed to range-balanced collimation model and LUCI SX pointing model from 2011 Jan 27. Re-executing acquire preset to M5_1133...

0327 - Nothing on the acquisition. Resetting IE/CA to some values Geno has in his notes from last night and doing a ptspiral to hunt. Andrew Rakich called in during this, and helped Geno through checking the pointing and collimation model setup. We got pointing corrections of
   IE = -75
   CA = +34
Andrew says these are close to what they had last night.

0340 - Active Preset BS9107 - failed. Clouds? Getting a heavy band over us again? No... the problem was that we had entered the M2 global offsets requested by Andrew into M1.

0349 - Corrected the global offsets, and resent the active preset to BS9107 - no luck, but the clouds do appear worse.

0354 - Acquire preset to ACT0280... got it? Taking an image through MODS1 red for additional confirmation. As we wait, the faint star we got is getting brighter - we have cloud problems...

Confirmed acquisition of ACT0280, went back to M5_1133, and reacquired it, so we have a decent pointing solution dialed in. However, relatively thick high clouds are rolling over us, so we're not having any luck getting fainter WFS stars to collimate yet. We'll wait out the clouds.

0407 - Wind has come up a little, with gusts up to 12m/s so Geno is closing the vent doors. Mirrors are close to ambient, so we're not too bad right now. Humidity is up to 12% from 4% earlier as well, temperature now about 4°C. Heavier cirrus overhead, but holes in the offing soon.

0440 - incoming big sucker hole... Going after BS9122 to collimate... Got a nice lock, collimation is converging. Seeing 0.8-1 arcsec and bouncing around a bit.

0500 - Well, had a little time. Got a nice lock with WFS rms around 150nm and at times 0.6-arcsec FWHM, but then we got hit with more clouds. During the cloud incursion, we saw a repeat of the problem from the September run where on an active preset it would grab the bright background at the upper left of the image as the "guide star" and go off into the weeds. The SExtractor parameter file and procedures for blocking overly-large FWHM "stars" from getting into the "brightest star" list has apparently not been implemented in the GCS. See
   /Repository/20110214/GCS/left_acquire000014.fits
   /Repository/20110214/GCS/left_acquire000015.fits
   /Repository/20110214/GCS/left_acquire000016.fits
for the failed acquisition sequence.

Checking the log (/lbt/log/2011/02/20110214.log), the FWHM of the brightest "star" it found was 264-277pixels (more than half the 470 pixel FoV of the guide camera).
   Feb 14 04:58:36 tcs4 LBT_left_GCS: brightest star @ 115.139999, 397.381989. FWHM = 264.329987, Ellipticity = 0.483000, Flux = 10451090.000000, Class = 0.000000
   Feb 14 04:58:44 tcs4 LBT_left_GCS: brightest star @ 117.107002, 399.071014. FWHM = 277.049988, Ellipticity = 0.492000, Flux = 10711810.000000, Class = 0.000000
   Feb 14 04:58:51 tcs4 LBT_left_GCS: brightest star @ 119.515999, 396.984009. FWHM = 271.750000, Ellipticity = 0.493000, Flux = 10761440.000000, Class = 0.000000
excerpting from the logs.

Of course the sucker hole closed about the same time...

0650 - Still waiting on the clouds to pass. There may be an opening in the next couple of hours...

1015 - ...or not. Some thinning, but the high clouds continue to cover most of the sky. The satellite is showing clearing ahead.

1120 - Pointing check WT10_294, moon has set, but thin high stuff still covers most of the sky. No problem acquiring, tweaking pointing
  IE = -80
  CA = +32

1125 - Collimation star BS9150. Started out fairly round and 1.2 arcsec seeing, before first WFS image taken. Converged at 0.9 arcsec FWHM

1143 - Move to Stone I field. Problem with guide star (bad typing in acq file - my bad). Converged, but we're facing into the wind so the seeing is a touch puffy. Mostly 1-arcsec FWHM

1146 - Integrating, had to abort as I'd forgotten to lock-on the IMCS (been up since early morning, getting tired). Getting guide warning about "GS elongated @ low WFS RMS". Images aren't great. Guide star is 12.5 with a nearby 17.2 companion which might also be contributing.

Instrument-Commanded Offset Tests

These tests are to verify propeof ther DETXY coordinate offsets and work on the crucial telescope interaction parts of the prototype long-slit target acquisition process.

1155 - Testing OFFSETXY (aka DETXY coords)
   offsetxy 10 0 rel
   offsetxy -20 0 rel
   offsetxy 10 -10 rel
   offsetxy  0 10 rel

1208 - Same test, now with field at PA=45°, different guide star. Same pattern. First image was before IMCS lock, repeated at end of sequence.

Worked great. Verified that the long x throw was indeed 20-arcsec using the measured red camera scale of 0.1228 arcsec/pix.

1220 - Setup to play with acquisition onto a long slit using the DETXY offsets. Slit: LS5x60x1.2 and only the red camera

Repeated for PA=0° and PA=45° A couple of different procedures (add together slit and sky images and make measures on the sum, measure individually). Worked very well.

Coordinates Signs:
    +/-X in DETXY moves the star +/-X on the CCD
    +/-Y in DETXY moves the star +/-Y on the CCD

1300 - Twilight coming up. We're done for the night.

-- RichardPogge - 14 Feb 2011
Topic revision: r6 - 16 Aug 2011, RichardPogge
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback